# SpineBench: A Clinically Salient, Level-Aware Benchmark Powered by the SpineMed-450k Corpus

### v1

@@ Abstract
[…]
@@ 2.2 Dataset Curation
To accurately extract comprehensive information from academic sources, we employed PaddleOCR (Du et al., 2020) to parse PDF documents and images from textbooks and literature. The output, containing both recognized text and layout analysis, was exported into Markdown format. This approach effectively preserved the structural integrity of the documents, including tables, figure placements, and overall layout. Furthermore, to ensure a precise mapping between figures, their captions, and corresponding contextual descriptions in the text, we developed a novel algorithm termed Picture Context Matching. The technical details of this algorithm are elaborated in the Appendix B.4.
@@
In close collaboration with medical experts, we designed a comprehensive annotation schema to generate high-quality, multi-task training data. The annotation process was tailored to the data source: (1) From External Knowledge Sources (e.g., Textbooks): We generated bilingual (Chinese and English) and multimodal (text and image-based) questions in both multiple-choice and open-ended formats using Expert VLM Model with carefully designed prompts. (2) From Opened-spine Datasets: We processed two open-source spinal datasets, Spark and Verse, to generate multi-turn question-and-answer dialogues that simulate doctor-patient interactions. These datasets consist mainly of unimodal 3D image slices (CT and MRI). To ensure consistency, we standardized the inputs by adaptively sampling 25 slices per case under clinical expert supervision. From this, we created over 300 simulated consultations to train models in their conversational abilities within spinal scenarios. (3) From Real Clinical Records: We created multiple-choice questions, multi-turn conversational datasets for patient interviews, and comprehensive spinal diagnostic reports via Expert VLM Model. For prompt design, please refer to the Appendix B.6
@@
A cornerstone of our dataset is the generation of detailed spinal diagnostic reports. In this process, we utilized real clinical reports from hospitals, incorporating physician recommendations, to design reports that encompass six dimensions, all aimed at simulating a complete clinical workflow: (1) Structured Imaging Findings: Analyze the provided medical images and distill key radiological evidence that supports the final diagnosis. (2) AI-Assisted Diagnosis: Formulate a diagnostic conclusion and articulate the reasoning process based on the synthesis of clinical data and imaging analysis. (3) Treatment Recommendations: This section is bifurcated to address different audiences. Patient-Centric Advice: Explain the rationale for the recommended surgical procedure in clear, non-technical language. Physician-Centric Rationale: Provide a robust, guideline-based decision tree to justify the surgical selection from a clinical perspective. (4) Risk and Prognosis Assessment: Conduct an objective evaluation of the potential risks and expected outcomes associated with the proposed surgical plan. (5) Postoperative Issue Management: Predict potential post-surgical complications for specific procedures and develop corresponding management strategies. (6) Diagnostic Rationale and Disclaimer: Provide complete diagnostic and surgical decision-making chain and disclaimer statement. Report examples are provided in the Appendix B.5.
@@ 5.1 Implementation Details
Finally, we conducted further training through multi-turn dialogues, report generation, and datasets comprising long-chain reasoning instructions. The goal of this stage is to develop the model’s advanced language comprehension and generation abilities, particularly in the contexts of dialogue interaction and report creation. All training details are provided in the Appendix B.2
@@ 5.2 Results on SpineBench
The evaluation results in Table 4 reveal severe limitations of current vision-language models (OpenAI, 2025a; Hurst et al., 2024; Google, 2025a) in medical domain applications. Large-scale open-source models perform particularly poorly: despite having 72B parameters, Qwen2.5-VL-72B (Bai et al., 2025) achieves only 79.88% average performance and a mere 63.80 cumulative score on medical report generation, far below practical application requirements. Even the best-performing open-source model GLM-4.5V (Hong et al., 2025) (83.26%) exhibits a nearly 6-point gap compared to the leading proprietary model Gemini-2.5-Pro (89.23%). This gap is more pronounced in medical report generation, where proprietary models exceed 85 points while open-source models struggle to reach 80. Additional medical report results are in the Appendix B.3.
@@ 6 CONCLUSIONS, LIMITATIONS, AND FUTURE WORK
7 Ethics Statement
@@ Appendix A LLM Usage
Appendix B Appendix
B.1 RELATED WORK
@@ B.2 Training Strategy
B.3 Performance comparison on medical report generation subtasks
@@
B.4 Picture Context Matching Algorithm
@@
B.5 QUANTITATIVE COMPARISON OF SpineGPT WITH GPT-4o

### v2

@@ Abstract
footnotetext: $\clubsuit$ Equal Contributions. 🖂Corresponding author: suntiansheng-@163.com; cfshan@nju.edu.cn
@@ 2.2 Dataset Curation
To accurately extract comprehensive information from academic sources, we employed PaddleOCR (Du et al., 2020) to parse PDF documents and images from textbooks and literature. The output, containing both recognized text and layout analysis, was exported into Markdown format. This approach effectively preserved the structural integrity of the documents, including tables, figure placements, and overall layout. Furthermore, to ensure a precise mapping between figures, their captions, and corresponding contextual descriptions in the text, we developed a novel algorithm termed Picture Context Matching. The technical details of this algorithm are elaborated in the Appendix F.
@@
In close collaboration with medical experts, we designed a comprehensive annotation schema to generate high-quality, multi-task training data. The annotation process was tailored to the data source: (1) From External Knowledge Sources (e.g., Textbooks): We generated bilingual (Chinese and English) and multimodal (text and image-based) questions in both multiple-choice and open-ended formats using Expert VLM Model with carefully designed prompts. (2) From Opened-spine Datasets: We processed two open-source spinal datasets, Spark and Verse, to generate multi-turn question-and-answer dialogues that simulate doctor-patient interactions. These datasets consist mainly of unimodal 3D image slices (CT and MRI). To ensure consistency, we standardized the inputs by adaptively sampling 25 slices per case under clinical expert supervision. From this, we created over 300 simulated consultations to train models in their conversational abilities within spinal scenarios. (3) From Real Clinical Records: We created multiple-choice questions, multi-turn conversational datasets for patient interviews, and comprehensive spinal diagnostic reports via Expert VLM Model. For prompt design, please refer to the Appendix H
@@
A cornerstone of our dataset is the generation of detailed spinal diagnostic reports. In this process, we utilized real clinical reports from hospitals, incorporating physician recommendations, to design reports that encompass six dimensions, all aimed at simulating a complete clinical workflow: (1) Structured Imaging Findings: Analyze the provided medical images and distill key radiological evidence that supports the final diagnosis. (2) AI-Assisted Diagnosis: Formulate a diagnostic conclusion and articulate the reasoning process based on the synthesis of clinical data and imaging analysis. (3) Treatment Recommendations: This section is bifurcated to address different audiences. Patient-Centric Advice: Explain the rationale for the recommended surgical procedure in clear, non-technical language. Physician-Centric Rationale: Provide a robust, guideline-based decision tree to justify the surgical selection from a clinical perspective. (4) Risk and Prognosis Assessment: Conduct an objective evaluation of the potential risks and expected outcomes associated with the proposed surgical plan. (5) Postoperative Issue Management: Predict potential post-surgical complications for specific procedures and develop corresponding management strategies. (6) Diagnostic Rationale and Disclaimer: Provide complete diagnostic and surgical decision-making chain and disclaimer statement. Report examples are provided in the Appendix G.
@@ 5.1 Implementation Details
Finally, we conducted further training through multi-turn dialogues, report generation, and datasets comprising long-chain reasoning instructions. The goal of this stage is to develop the model’s advanced language comprehension and generation abilities, particularly in the contexts of dialogue interaction and report creation. All training details are provided in the Appendix D
@@ 5.2 Results on SpineBench
The evaluation results in Table 4 reveal severe limitations of current vision-language models (OpenAI, 2025a; Hurst et al., 2024; Google, 2025a) in medical domain applications. Large-scale open-source models perform particularly poorly: despite having 72B parameters, Qwen2.5-VL-72B (Bai et al., 2025) achieves only 79.88% average performance and a mere 63.80 cumulative score on medical report generation, far below practical application requirements. Even the best-performing open-source model GLM-4.5V (Hong et al., 2025) (83.26%) exhibits a nearly 6-point gap compared to the leading proprietary model Gemini-2.5-Pro (89.23%). This gap is more pronounced in medical report generation, where proprietary models exceed 85 points while open-source models struggle to reach 80. Additional medical report results are in the Appendix E.
@@ 6 CONCLUSIONS, LIMITATIONS, AND FUTURE WORK
APPENDIX
Contents
A  Checklist ........................................................................................................................................................................A
A.1 Ethics Statement ........................................................................................................................................................................A.1
A.2 Reproducibility Statement ........................................................................................................................................................................A.2
A.3 LLM Usage ........................................................................................................................................................................A.3
B  Contributions and Acknowledgments ........................................................................................................................................................................B
C  Related Work ........................................................................................................................................................................C
D  Training Strategy ........................................................................................................................................................................D
E  Performance Comparison ........................................................................................................................................................................E
F  Picture Context Matching Algorithm ........................................................................................................................................................................F
G  Quantitative Comparison of SpineGPT with GPT-4O ........................................................................................................................................................................G
H  Prompts ........................................................................................................................................................................H
Appendix A Checklist
A.1 Ethics Statement
@@ A.3 LLM Usage
Appendix B Contributions and Acknowledgments
  1. 1.
    Ming Zhao
    $\pi^{3}$ Lab, Jilin University
  2. 2.
    Wenhui Dong
    $\pi^{3}$ Lab, Nanjing University
  3. 3.
    Yang Zhang
    The Fourth Medical Center of People’s Liberation Army General Hospital
  4. 4.
    Xiang Zheng
    $\pi^{3}$ Lab, Institute of Automation, Chinese Academy of Sciences
  5. 5.
    Zhonghao Zhang
    $\pi^{3}$ Lab, Ningxia University
  6. 6.
    Zi An Zhou
    $\pi^{3}$ Lab, Zhejiang University
  7. 7.
    Yunzhi Guan
    Huashan Hospital, Fudan University
  8. 8.
    Liukun Xu
    The Fourth Medical Center of People’s Liberation Army General Hospital
  9. 9.
    Wei Peng
    Stanford University
  10. 10.
    Zhaoyang Gong
    Huashan Hospital, Fudan University
  11. 11.
    Zhicheng Zhang
    The Fourth Medical Center of People’s Liberation Army General Hospital
  12. 12.
    Dachuan Li
    Huashan Hospital, Fudan University
  13. 13.
    Xiaosheng Ma
    Huashan Hospital, Fudan University
  14. 14.
    Yuli Ma
    Sanyou Medical
  15. 15.
    Jianing Ni
    $\pi^{3}$ Lab
  16. 16.
    Changjiang Jiang
    Wuhan University
  17. 17.
    Lixia Tian
    Beijing Jiaotong University
  18. 18.
    Qixin Chen
    The Second Affiliated Hospital, Zhejiang University
  19. 19.
    Kaishun Xia
    The Second Affiliated Hospital, Zhejiang University
  20. 20.
    Pingping Liu
    Jilin University
  21. 21.
    Tongshun Zhang
    Jilin University
  22. 22.
    Zhiqiang Liu
    $\pi^{3}$ Lab,Huazhong University of Science and Technology
  23. 23.
    Zhongan Bi
    $\pi^{3}$ Lab,Zhejiang University
  24. 24.
    Chenyang Si
    Nanjing University
  25. 25.
    Tiansheng Sun
    The Fourth Medical Center of People’s Liberation Army General Hospital
  26. 26.
    Caifeng Shan
    Nanjing University
Appendix C RELATED WORK
@@ Appendix D Training Strategy
Appendix E Performance comparison on medical report generation subtasks
@@
Appendix F Picture Context Matching Algorithm
@@
Appendix G QUANTITATIVE COMPARISON OF SpineGPT WITH GPT-4o